In a right circular cone, the radius of the base is half as long as the altitude. If an error of 2% is made in measuring the radius, find the percentage of error made in the computed volume.

1
Expert's answer
2021-11-22T18:15:07-0500

Let "h" be the altitude and "r" be the radius of the base of a right circular cone.

Given: "h=2r"  and "\Delta r/r=0.02." inches.

We know that the volume of the right circular is "V=\dfrac{1}{3}\pi r^2h."

Now "h=2r\Rightarrow V=\dfrac{1}{3}\pi r^2(2r)=\dfrac{2}{3}\pi r^3."

Take the differentials on both sides


"dV=2\pi r^2dr"

Then


"\dfrac{dV}{V}=\dfrac{2\pi r^2dr}{\dfrac{2}{3}\pi r^3}=3\dfrac{dr}{r}=3(0.02)=0.06"

The percentage of error made in the computed volume is 6%.